# Value Alignment of Social Media Ranking Algorithms

### v1

@@ 3.1.2. Validating classifier performance
To validate the performance of the LLM classifier, we compare the generated labels with a dataset of 4,562 Twitter/X posts, where each post is labeled by $\geq 4$ human annotators. For each post, we report LLM-Consensus MAE, or the mean absolute error (MAE) between the LLM-generated label and the consensus label (e.g., mean human rating). We then compare this error to the mean disagreement that an individual human annotator has with the consensus label (Human-Consensus MAE). To formalize this, for each of the $n$ annotators who labeled a value $v$ on a given post, we compute the MAE between the label they assigned to the value and the mean labels of the remaining n1n-1 annotators as follows: $\frac{1}{n}\sum_{i=0}^{n}|v_{i}-\frac{1}{n-1}\sum_{j=0,j\neq i}^{n}v_{j}|$.
Across all values the LLM-Consensus MAE ($0.95\pm 1.10$) is lower than the Human-Consensus MAE ($1.07\pm 1.05$). This result suggests that the LLM-generated labels agree with the consensus more than an individual human annotator agrees with the consensus. For all but three values (“Stimulation,” “Hedonism”, and “Achievement”), the LLM performs as well as, if not better than, human annotators when it comes to predicting the consensus value label. For the three values where Human-Consensus MAE is less than the LLM-Consensus MAE, the difference between the two is at most $0.21$. See Appendix for results disaggregated by value.
@@ 3.2. Integrating multiple values together to rank content
For example, suppose that one user, Jeanne, wants to rank their feed by Self-Transcendence values (e.g., “Caring”, “Dependability”, “Universal Concern”); another user, Jeff, wants to rank their feed by Self-Enhancement values that stand opposite of Self-Transcendence values (e.g., “Achievement”, “Dominance”, “Resources”) in Schwartz’s circumplex. They each assign a weight vector $w$ with a weight of $1$ to the values they prioritize, 1-1 to the values on the opposite side of the circumplex, and leave all others at the default 0 weight. Then, a tweet expressing strong support for “Caring” might get a score of $6$ from the classifier; ranking it highly for Jeanne as its score is $6\times 1=6$, but at the bottom of the feed for Jeff, whose score is $6\times-1=-6$.
@@
Our approach is an example of how to navigate value elicitation and trade-offs, but is not the only way. Prior work has articulated other approaches such as reward learning from human feedback and metric elicitation, which offer algorithmic alternatives (Christiano et al., 2017; Hiranandani et al., 2019). An important finding from this work is that more complicated trade-offs between values that may be non-linear can still be expressed using simpler linear trade-offs (Hiranandani et al., 2019; Laidlaw and Russell, 2021; Robertson et al., 2023). This equivalence indicates that linear approaches such as ours are robust enough to capture even more complex trade-offs that occur.
@@ 3.3. Obtaining users’ value preferences
The final part of our pipeline is eliciting what values users may want to see (or not see) in their feeds: the weight vector $w$ on values. There is no single strategy for collecting these preferences. One method can be to learn value preferences from users’ past behavior. Alternatively, users can directly input what values the want to see. Within this domain of end-user control, there are many options for how users can specify their values — from having expert interfaces to simpler but less expressive controls — that each come with their own trade-offs between afforded control and potential effort or cognitive burden. Our Study 1 below utilizes a heavyweight method of delivering a survey instrument to measure users’ value priorities, then ranking feeds using the user’s top values from their survey. However, it is possible that what users want to see in their feed may not align with their personal values. In Study 2, we offer users a set of sliders corresponding to each of Schwartz’s values, offering more fine-grained control. Slider weights range from -1 to 1 in increments of 0.25. Since users may be unfamiliar with the values, we provide a tooltip next to each slider with the definition.
@@ 4.1.1. Study Design
In the first step, to collect tweets from the participants’ Twitter feed, we ask them to download a browser extension that we have developed. The extension makes requests to the Twitter API on behalf of the user for fetching several consecutive batches of tweets. A “batch” is defined as the set of tweets returned from one call to the Twitter API, excluding those tweets that are “promoted” (advertisements) and recommendations about who to follow or to subscribe to. We collect our batches from the user’s “For You” timeline (the engagement feed). By using tweets from the engagement feed, we can control for user interest and satisfaction with the content. It also provides a more realistic window into how social media companies may adopt value-based ranking mechanisms on top of existing algorithms. These tweets are the inventory of content being ranked on our platform.
Participants are then redirected to our external experimental platform website, where they complete the Portrait Value Questionnaire (PVQ) (Schwartz, 2012b), a 57-item survey used to measure an individual’s prioritization of the 19 Schwartz values. We use the results from this survey to determine the weights for the value-aligned feed. In each trial for Study 1, we test how well our approach can rank by a single value at a time. To do so, we set the weight of one value equal to 1, and the weight of all other values to 0. Specifically, for the first trial, we select only the participant’s highest-ranked value from the PVQ, subsetting to only the values for which there is at least one tweet in the inventory expressing this value. Next, in the second trial, we look at the second highest-ranked value that also appears in our inventory, and so on. Each participant thus sees a set of values that reflect their top-ranked values as rated in the PVQ. There are four such trials, each focused on a different highly-rated value for that participant.
@@ 4.1.2. Participants
We recruited participants in March 2025 via the Prolific platform. To qualify for our studies, participants must reside in the United States, be over 18 years old, have a Twitter account, and pass quality checks on Prolific (i.e., over 250 prior submissions and $\geq 95$% acceptance rate). Since values may differ depending on individuals’ political orientation  (Schwartz, 2012a; Haidt and Joseph, 2004; Graham et al., 2009), we sampled participants from across the political spectrum to ensure we are not systematically excluding any values. We launched separate tasks on Prolific to recruit participants who self-identified as Democrats and for those who self-identified as either Republican or Independent. [^6] We combined recruitment for Republicans and Independents to ensure adequate sample size. Participants received \$4.25 for completing the task, and those that correctly identified all four value re-ranked feeds received a bonus of $\mathdollar 0.50$. We determined the payment from pilot studies and our intention to provide a rate of \$15 per hour. In total, we collected responses from 194 participants and removed those that failed our attention checks ($N=17$) or provided incomplete responses ($N=36$). For our analyses, we were left with 564 responses from 141 participants.
@@ 4.2.2. Most—but not all—values produce recognizable feeds
We also conducted a disaggregated analysis to examine recognizability across the 19 values. We find that for all but one of the values (“Interpersonal Conformity”), mean recognizability is above $50\%$ (Table 2). We also compute $95\%$ confidence intervals by bootstrapping over 1,000 iterations and find that all but three (“Interpersonal Conformity”, “Resources”, “Stimulation”) have recognizability rates significantly higher than random chance. We observe a wide range of recognizability, spanning from 100.0100.0% for “Preservation of Nature” to $47.1\%$ for “Interpersonal Conformity”. Values grouped under the category Openness to Change (e.g., “Self-directed actions”, “Stimulation”) had relatively lower recognizability, with a mean of $67.6\%$, compared to the Self-Transcendence group (e.g., “Caring”, “Humility”), which has a mean recognizability of $86.1\%$. One explanation for this variance may be the interpretability of different values. Even though we translated Schwartz’s constructs into more understandable terms, certain constructs are easier to convey than others. For example, understanding what “Caring” means, which scores higher on recognizability ($89.7\%$), may be easier for participants who are unfamiliar with Schwartz’s Basic Values compared to a construct such as “Self-directed actions” ($62.0\%$). Another factor may be because the underlying engagement ranking already filters content, implicitly selecting for certain values, such as those related to Openness to Change. As a result, feeds ranked by these values may have produced less pronounced differences.
@@ 5.1. Method
We conducted a between-subjects experiment (N=250) eliciting participants’ value weights using a set of sliders (see Fig. 5). Participants were assigned into one of six conditions, corresponding to the number of value sliders that participants will be able to adjust. The number of value sliders shown ranges from 1 to 5 or all 19 value sliders (i.e., “Full” condition).
@@ 5.1.1. Study Design
In the experimental task, the user is exposed to a “training” and “testing” phases. During the training phase, the participant uses the controls to create a feed that best matches their values. Depending on the condition, participants are able to change only a select number of value sliders (or all if they are in the Full condition). They must change at least one value to proceed. After the training phase, the user answers a survey on the perceived complexity of the interface as well as providing a free-text response on what they liked and did not like about the interface. In the testing phase, similar to Study 1, users are shown a side-by-side comparison of their value-aligned feed and engagement feed (i.e., the feed as ranked by Twitter/X’s algorithm), without labels and in a randomized position. They are asked to identify which feed has been ranked using the weights from their input across four trials. In each trial, the tweets shown to the participant come from a different set of batches. For each trial, the pool of possible tweets being ranked consists of up to seven batches (roughly 210 tweets), and each feed shows only the top 20 ranked tweets by its own ranking function.
@@ 5.1.3. Measures
We focus on two main measures in Study 2: recognizability and complexity. Similar to Study 1, recognizability is measured as the percentage of trials in which participants correctly select the value-aligned feed in an unlabeled side-by-side comparison (Fig. 4). In Study 2, we add an additional dependent variable, complexity. To measure complexity, we adapt questions from the NASA Task Load Index (TLX) (Hart, 1988), participants rated how hard (“I had to work hard to create the resulting feed.”) and mentally demanding (“Using the controls was mentally demanding.”) each interface was to use on a 7-point Likert scale (1=“Strongly disagree”, 7=“Strongly agree”). We average over the responses as our measure of complexity (Cronbach’s $\alpha=0.66$). The survey measures are gathered after the training round but before the testing rounds.
@@ 5.2.1. Value aligned feeds become harder to recognize as more values are activated
Overall, we observe a weak negative association between the number of sliders changed and recognizability ($\beta=-0.028$), but this effect is not statistically significant ($p=0.14$). As shown in Fig. 6, a closer inspection reveals a noticeable drop in recognizability when comparing participants who changed exactly one value ($72.6\%\pm 30.1$) to those who changed more than one ($61.5\%\pm 29.1$). To formalize this observation, we conduct an exploratory analysis, replacing Sliders Changed in our linear mixed-effect model with a binarized predictor >1 Slider Changed. >1 Slider Changed is equal to 1 when the number of sliders changed is greater than 1 and 0 otherwise. We observe a significant negative correlation between >1 Slider Changed and recognizability ($\beta=-0.60$, $p=0.02$). In short, recognizing value aligned feeds becomes more difficult when multiple values are activated, likely because participants are forced to reason over more concepts each exerting different degrees of influence on the feed, increasing their cognitive load. However, beyond the initial drop-off from re-ranking using a single to multiple values, recognizability does not decline significantly as the number of value dimensions increases.
5.2.2. Complexity does not increase across conditions
We also examined whether participants perceived the interface as being more complex when they are able to adjust more sliders. Participants in the single-slider condition reported the lowest perceived complexity (M = $1.42$, SD = $1.27$) compared to conditions where they could adjust multiple sliders. However, independent $t$-tests comparing all six conditions, with Bonferroni correction applied to control for multiple comparisons, revealed no statistically significant differences in reported complexity. Qualitative feedback provides additional insight into this pattern. Many participants described the cognitive effort involved in prioritizing which values to adjust, noting that it was challenging to decide which sliders mattered most (e.g., “it was difficult to have to choose only two max that I could adjust”) or expressed frustration about needing to limit their changes (“there were a lot of sliders that I felt were important to me, so it was tough only choosing five”). Conversely, some participants in the Full Sliders condition reported experiencing choice overload, remarking that there “were too many options to choose from”, leading them to feel overwhelmed by choice fatigue.
@@ 6. Understanding Users’ Value Selections
Finally, we investigate how and when users want to use value-based controls to curate their feeds. While our earlier studies examined whether value-based re-ranking can produce recognizable feeds, here we focus on how participants make sense of these controls and in what contexts they find them useful. We adopt a mixed-methods approach, combining qualitative insights from a think-aloud study with quantitative analyses of participants’ value selections and the resulting re-ranked feeds.
6.1. Method
We conduct a mixed-methods study, combining qualitative data from a think-aloud study where participants interact with the interface from Study 2 as with a quantitative analysis of the value selections and resulting feeds. For the qualitative think-aloud study, we recruited thirteen participants via university mailing lists. During the study, participants completed the same protocol as described in Study 2 (Sec. 5) and were given access to all 19 sliders (i.e., the Full Sliders condition). We instructed the participants to think aloud as they adjusted their value sliders and reason about which feed they believed was value-ranked. Each session lasted approximately 35–60 minutes, and participants received a \$20 gift card via Tremendous as compensation. We pair this qualitative data with a quantitative analysis of the slider selections and resulting feeds for the 47 participants from Study 2 who were assigned to the Full Sliders condition.
6.2. Results
6.2.1. What values do users select to re-rank their feeds?

### v2

@@ 3.1.2. Validating classifier performance
To validate the performance of the LLM classifier, we compare the generated labels with a dataset of 4,562 Twitter/X posts from Epstein et al. (2025), in which each post is labeled by $\geq 4$ human annotators. For each post, we report LLM-Consensus MAE, or the mean absolute error (MAE) between the LLM-generated label and the consensus label (e.g., mean human rating). We then compare this error to the mean disagreement that an individual human annotator has with the consensus label (Human-Consensus MAE). To formalize this, for each of the $n$ annotators who labeled a value $v$ on a given post, we compute the MAE between the label they assigned to the value and the mean labels of the remaining n1n-1 annotators as follows: $\frac{1}{n}\sum_{i=0}^{n}|v_{i}-\frac{1}{n-1}\sum_{j=0,j\neq i}^{n}v_{j}|$.
Across all values the LLM-Consensus MAE ($0.95\pm 1.10$) is lower than the Human-Consensus MAE ($1.07\pm 1.05$). This result suggests that the LLM-generated labels agree with the consensus more than an individual human annotator agrees with the consensus. For all but three values (“Stimulation,” “Hedonism”, and “Achievement”), the LLM performs as well as, if not better than, human annotators when it comes to predicting the consensus value label. For the three values where Human-Consensus MAE is less than the LLM-Consensus MAE, the difference between the two is at most $0.21$. See Appendix D.1 for results disaggregated by value and additional evaluation metrics.
@@ 3.2. Integrating multiple values together to rank content
Side-by-side comparison of two simulated feeds showing the effect of different value weights. Side-by-side comparison of two simulated feeds showing the effect of different value weights. The left feed prioritizes “Achievement” and “Dominance” while downranking “Humility,” whereas the right feed prioritizes “Humility” while downranking “Face” (reputation), illustrating how ranking changes based on user-defined values
@@
For example, suppose that one user, Jeanne, wants to rank their feed by Self-Transcendence values (e.g., “Caring”, “Dependability”, and “Universal Concern”); another user, Jeff, wants to rank their feed by Self-Enhancement values that stand opposite of Self-Transcendence values (e.g., “Achievement”, “Dominance”, and “Resources”) in Schwartz’s circumplex. They each assign a weight vector $w$ with a weight of $1$ to the values they prioritize, 1-1 to the values on the opposite side of the circumplex, and leave all others at the default 0 weight. Then, a tweet expressing strong support for “Caring” might get a score of $6$ from the classifier; ranking it highly for Jeanne as its score is $6\times 1=6$, but at the bottom of the feed for Jeff, whose score is $6\times-1=-6$.
@@
Our approach is an example of how to navigate value elicitation and trade-offs, but is not the only way. Prior work has articulated other approaches such as reward learning from human feedback and metric elicitation, which offer algorithmic alternatives (Christiano et al., 2017; Hiranandani et al., 2019). In contrast, we adopt a linear weighting scheme over values, which is a common choice in industry ranking pipelines (Milli et al., 2023; Smith, 2021; Twitter Engineering, 2023). While linear weights cannot capture all potential interactions between values — for example, preferences that depend on specific combinations of values — they provide an interpretable starting point for enabling value-based ranking.
@@ 3.3. Obtaining users’ value preferences
The final part of our pipeline is eliciting what values users may want to see (or not see) in their feeds: the weight vector $w$ on values. There is no single strategy for collecting these preferences. One method can be to learn value preferences from users’ past behavior. Alternatively, users can directly input what values they want to see. Within this domain of end-user control, there are many options for how users can specify their values — from having expert interfaces to simpler but less expressive controls — that each come with their own trade-offs between afforded control and potential effort or cognitive burden. Our Study 1 below utilizes a heavyweight method of delivering a survey instrument to measure users’ value priorities, then ranking feeds using the user’s top values from their survey. However, it is possible that what users want to see in their feed may not align with their personal values. In Study 2, we offer users a set of sliders corresponding to each of Schwartz’s values, offering more fine-grained control. Slider weights range from -1 to 1 in increments of 0.25. Since users may be unfamiliar with the values, we provide a tooltip next to each slider with the definition.
@@ 4.1. Method
Screenshot of the experimental platform interface. Screenshot of the experimental platform interface. Two unlabeled feeds are displayed side by side (“Feed A” and “Feed B”), visually replicating Twitter’s timeline style, used to test participants’ ability to distinguish between engagement-driven and value-aligned rankings
@@ 4.1.1. Study Design
In the first step, to collect tweets from the participants’ Twitter feed, we ask them to download a browser extension that we have developed. The extension makes requests to the Twitter API on behalf of the user for fetching several consecutive batches of tweets. A “batch” is defined as the set of tweets returned from one call to the Twitter API, excluding those tweets that are “promoted” (advertisements) and recommendations about who to follow or to subscribe to. We collect our batches from the user’s “For You” timeline (the engagement feed). By using tweets from the engagement feed, we can control for user interest and satisfaction with the content. It also provides a more realistic window into how social media companies may adopt value-based ranking mechanisms on top of existing algorithms. These tweets are the inventory of content being ranked on our platform. Although the content on users’ “For You” timelines varies, leading to differences in the posts that participants see, we chose to use posts from participants’ own feeds rather than a static set across users, as this setup more realistically reflects real-world feed ranking deployment.
Participants are then redirected to our external experimental platform website, where they complete the Portrait Value Questionnaire (PVQ) (Schwartz, 2012), a 57-item survey used to measure an individual’s prioritization of the 19 Schwartz values. We use the results from this survey to determine the weights for the value-aligned feed. In each trial for Study 1, we test how well our approach can rank by a single value at a time. To do so, we set the weight of one value equal to 1, and the weight of all other values to 0. Specifically, for the first trial, we select only the participant’s highest-ranked value from the PVQ, subsetting to only the values for which there is at least one tweet in the inventory expressing this value. Next, in the second trial, we look at the second highest-ranked value that also appears in our inventory, and so on. Each participant thus sees a set of values that reflect their top-ranked values as rated in the PVQ. There are four such trials, each focused on a different highly-rated value for that participant.
@@ 4.1.2. Participants
We recruited participants in March 2025 via the Prolific platform. To qualify for our studies, participants must reside in the United States, be over 18 years old, have a Twitter account, and pass quality checks on Prolific (i.e., over 250 prior submissions and $\geq 95$% acceptance rate). Since values may differ depending on individuals’ political orientation  (Schwartz, 2011; Haidt and Joseph, 2004; Graham et al., 2009), we sampled participants from across the political spectrum to ensure we are not systematically excluding any values. We launched separate tasks on Prolific to recruit participants who self-identified as Democrats and for those who self-identified as either Republican or Independent. [^7] We combined recruitment for Republicans and Independents to ensure adequate sample size. Participants received \$4.25 for completing the task, and those that correctly identified all four value re-ranked feeds received a bonus of $\mathdollar 0.50$. We determined the payment from pilot studies and our intention to provide a rate of \$15 per hour. In total, we collected responses from 194 participants and removed those that failed our attention checks ($N=17$) or provided incomplete responses ($N=36$). For our analyses, we were left with 564 responses from 141 participants.
@@ 4.2.2. Most — but not all — values produce recognizable feeds
We also conducted a disaggregated analysis to examine recognizability across the 19 values. We find that for all but one of the values (“Interpersonal Conformity”), mean recognizability is above $50\%$ (Table 2). We also compute $95\%$ confidence intervals by bootstrapping over 1,000 iterations and find that all but three (“Interpersonal Conformity”, “Resources”, and “Stimulation”) have recognizability rates significantly higher than random chance. We observe a wide range of recognizability, spanning from 100.0100.0% for “Preservation of Nature” to $47.1\%$ for “Interpersonal Conformity”. Values grouped under the category Openness to Change (e.g., “Self-directed actions” and “Stimulation”) had relatively lower recognizability, with a mean of $67.6\%$, compared to the Self-Transcendence group (e.g., “Caring” and “Humility”), which has a mean recognizability of $86.1\%$. One explanation for this variance may be the interpretability of different values. Even though we translated Schwartz’s constructs into more understandable terms, certain constructs are easier to convey than others. For example, understanding what “Caring” means, which scores higher on recognizability ($89.7\%$), may be easier for participants who are unfamiliar with Schwartz’s Basic Values compared to a construct such as “Self-directed actions” ($62.0\%$). Another factor may be because the underlying engagement ranking already filters content, implicitly selecting for certain values, such as those related to Openness to Change. As a result, feeds ranked by these values may have produced less pronounced differences.
@@ 5. Study 2: Re-ranking by Multiple, User-Controlled Values
Screenshot of the interface for Study 2 Screenshot of the interface for Study 2, where participants could adjust slider weights for up to 19 values. The left side shows the feed being re-ranked in real time, while the right side presents the 19 sliders.
@@ 5.1. Method
We conducted a between-subjects experiment (N=250) eliciting participants’ value weights using a set of sliders (see Fig. 5). Participants were assigned into one of six conditions. While participants were shown sliders corresponding to all nineteen values, their assigned conditioned determined the maximum number of sliders they could adjust. The number of value sliders participants could adjust ranged from 1 to 5 or all 19 value sliders (i.e., “Full” condition). Our pilot tests, described in more detail in Appendix B, included conditions for adjusting 1 to 6 values as well as the full 19. The results suggested that recognizability began to plateau beyond just a few values and informed our pre-registered final study, where we included 1 to 5 sliders and all 19.
@@ 5.1.1. Study Design
In the experimental task, the user is exposed to a “training” and “testing” phases. During the training phase, the participant uses the controls to create a feed that best matches their values. Depending on the condition, participants are able to change only a select number of value sliders (or all if they are in the Full condition). They must change at least one value to proceed. After the training phase, the user answers a survey on the perceived complexity of the interface as well as providing a free-text response on what they liked and did not like about the interface. In the testing phase, similar to Study 1, users are shown a side-by-side comparison of their value-aligned feed and engagement feed (i.e., the feed as ranked by Twitter/X’s algorithm), without labels and in a randomized position. As a robustness check, we replicated the conditions in which participants could adjust either a single slider or all 19, but instead asked them to choose between the value-aligned feed and a randomly reshuffled feed. We found qualitatively similar results (see Appendix D.2.1 for full details). They are asked to identify which feed has been ranked using the weights from their input across four trials. In each trial, the tweets shown to the participant come from a different set of batches. For each trial, the pool of possible tweets being ranked consists of up to seven batches (roughly 210 tweets), and each feed shows only the top 20 ranked tweets by its own ranking function.
@@ 5.1.3. Measures
We focus on two main measures in Study 2: recognizability and task load. Similar to Study 1, recognizability is measured as the percentage of trials in which participants correctly select the value-aligned feed in an unlabeled side-by-side comparison (Fig. 4). In Study 2, we add an additional dependent variable, task load. To measure task load, we adapt questions from the NASA Task Load Index (TLX) (Hart, 1988), participants rated how hard (“I had to work hard to create the resulting feed.”) and mentally demanding (“Using the controls was mentally demanding.”) each interface was to use on a 7-point Likert scale (1=“Strongly disagree”, 7=“Strongly agree”). We average over the responses as our measure of task load (Cronbach’s $\alpha=0.66$). The survey measures are gathered after the training round but before the testing rounds.
@@ 5.2. Results
Bar chart showing recognizability rates by number of sliders adjusted in Study 2. Bar chart showing recognizability rates by number of sliders adjusted in Study 2. Recognizability declines as more values are simultaneously activated, from 72.6% for single-value ranking to 64.0% for 2 sliders, 65.5% for 3 sliders, 59.3% for 4 sliders, 57.9% for 5 sliders, and 61.0% for 5+ sliders.
@@ 5.2.1. Value aligned feeds become harder to recognize as more values are activated
Overall, we observe a weak negative association between the number of sliders changed and recognizability ($\beta=-0.028$), but this effect is not statistically significant ($p=0.14$). As shown in Fig. 6, a closer inspection reveals a noticeable drop in recognizability when comparing participants who changed exactly one value ($72.6\%\pm 30.1$) to those who changed more than one ($61.5\%\pm 29.1$). To formalize this observation, we conduct an exploratory analysis, replacing Sliders Changed in our linear mixed-effect model with a binarized predictor >1 Slider Changed. >1 Slider Changed is equal to 1 when the number of sliders changed is greater than 1 and 0 otherwise. We observe a significant negative correlation between >1 Slider Changed and recognizability ($\beta=-0.60$, $p=0.02$). In short, recognizing value aligned feeds becomes more difficult when multiple values are activated, likely because participants are forced to reason over more concepts each exerting different degrees of influence on the feed, increasing their cognitive load (Popowski et al., 2026). However, beyond the initial drop-off from re-ranking using a single to multiple values, recognizability does not decline significantly as the number of value dimensions increases.
5.2.2. Task load does not increase across conditions
We also examined whether participants perceived the interface as being more complex when they are able to adjust more sliders. Participants in the single-slider condition reported the lowest perceived task load (M = $1.42$, SD = $1.27$) compared to conditions where they could adjust multiple sliders. However, independent $t$-tests comparing all six conditions, with Bonferroni correction applied to control for multiple comparisons, revealed no statistically significant differences in reported task load. Qualitative feedback provides additional insight into this pattern. Many participants described the cognitive effort involved in prioritizing which values to adjust, noting that it was challenging to decide which sliders mattered most (e.g., “it was difficult to have to choose only two max that I could adjust”) or expressed frustration about needing to limit their changes (“there were a lot of sliders that I felt were important to me, so it was tough only choosing five”). Conversely, some participants in the Full Sliders condition reported experiencing choice overload, remarking that there “were too many options to choose from”, leading them to feel overwhelmed by choice fatigue.
@@ 6. Understanding Users’ Value Selections
Finally, we investigate how and when users want to use value-based controls to curate their feeds. While our earlier studies examined whether value-based re-ranking can produce recognizable feeds, here we focus on how participants make sense of these controls and in what contexts they find them useful. We conduct a quantitative analysis of the value selections and resulting feeds for the 47 participants from Study 2 who were assigned to the Full Sliders condition.
6.1. What values do users select to re-rank their feeds?